OpenStudy (anonymous):

Multiplication Property of Negative One Prove that -1x X= -X a.−1×x+x=−1×x+1×x b.=(−1+1)(x) c.=0×x d. =0 e. =−x+x f.∴−1×x+x=−x+x g.∴−1×x=−x could anyone help me w/ the reasons for this??? like mult. prop. of = or sth...

OpenStudy (anonymous):

HELP!!!

OpenStudy (anonymous):

HELP!!!

OpenStudy (anonymous):

HELP!!!

OpenStudy (anonymous):

(-)*(+)=(-) (-)*(-)=(+) (+)*(+)=(+)

OpenStudy (anonymous):

wat?????

OpenStudy (anonymous):

I mean like as in reasons... Those are the thingies on the left side... on a proof...